# Super-Resolution Generative Adversarial Networks based Video Enhancement Kağan ÇETİN

### v1

@@ 1.1 Background
Figure 1: Basic overview of a GAN architecture
@@
Figure 2: Basic overview of an SRGAN architecture
@@
Also in the following year in 2018, another architecture called as ESRGAN is being publicated and shown the literature that further upgrades can still be made on residual blocks by stacking them consecutively and let them share information among each other. Each output of those blocks are stored and shared directly with the any other residual block waiting on the flow way. The desired architecture is shown below in Fig. (5) and called as Residual in Residual Dense Block (RRDB)  [15].
Figure 5: Residual in Residual Dense Block
@@
Figure 6: a. Leaky ReLU compared to b. (ordinary) ReLU
@@
Figure 7: Spacetime View of Non-Local Block
@@ 2.1 Spatio-Temporal Learning Architecture built for Video Super-Resolution
At first, it should be noted that two different experimental and basic models for different complexities of jobs are designed. The first shown in Figure 9 is built using consecutive RRDBs mainly, whereas the other model which is shown in Figure 9 is built based on successive residual blocks. Both of the models can be split into 4 different parts, that they can be called channel expanding and temporal feature extraction layer, spatial feature extraction layer, reconstruction layer and upsampling and channel compression layer. The RRDBs are based on Residual Blocks, and each residual block is simply a sequentially connected series of convolutions with skip connections, where flow diagram of those blocks are shown in Figure 10.a and shown in Figure 10.b. In simple terms, each RRDB is also a series (3) of residual blocks with another skip connection at the end.
@@
Figure 10: a. Residual in Residual Dense Block and b. Residual Block modules designed for this project.
@@
Assuming $\sigma(y)=LeakyReLU(y)$, a convolutional operation followed by a leaky relu function in the flow diagram mathematically represents the following Equation 5.
@@
In the end, output of the residual block used in this project is as in given in Equation 6 below.
@@
Note the following notations represent the same thing, given in Equation 7 below.
@@
Correspondingly, RRDB class designed for this project is as follows given in Equation 8 below.
@@
On the other hand, Non-Local Blocks has a different calculation logic from which its spatial feature extraction ability comes from. Normally, ordinary 2D convolution blocks and its derivations are based on extracting the temporal relationship, which limits the model’s learning ability to 2D dataset only. However, with non-local blocks, even in its simplest form, which is called dot product based non-local blocks, we get the ability to learn the 3D relationships by calculating the weighted sum of the 3D input sets. Mathematically, it is represented in a very generic form given in Equation 9  [5].
@@
where the output of a non-local block indeed is as follows in Equation 10
@@
  • $g(\cdot)$: It is the unary function g (given in Equation 12) computes a representation of the input signal at the position $j$.
@@
The normalization factor $C(\cdot)$ is given below in Equation 11.
@@
The unary function $g(\cdot)$ is given below in Equation 9.
@@
Pairwise function $f(\cdot)$ used in this function (dot product) is given below in Equation 13.
@@
    $f(x_{i},x_{j})=\sigma(w_{f}^{T}\cdot\theta(x_{i})^{T}\mathbin{\|}\phi(x_{j})),\sigma(\cdot)=ReLU(\cdot)$ (19)
Figure 11: Non-Local Block module used in this project.
Figure 12: U-Net architecture used as discriminator
In this adaptation process, the purpose of the discriminator shall not be changed deeply. Since their performance on image segmentation is proven and it is shown that their encoder & decoder based architecture is reliably well-performing along large scales of projects  [34], a U-Net architecture (see Figure 12) is built as the discriminator and it can be kept the same as long as it does well in distinguishing artificially made images from real images at first. After a while, it will converge to a non-zero fixed amount of error. Attempting to minimize this error excessively, bring it very close to zero, and significantly increase the complexity of the discriminator block in the process will negatively affect the overall procedure. The backpropagation of a non-zero loss from the discriminator block to the generator is necessary for a healthy process.
@@ 3.1.7 Evaluation and Definition of Loss Functions
    Figure 13: a. Moore Neighborhood compared to b. Von Neumann Neighborhood
@@
    Figure 14: Evaluation of Laplacian Pyramid
@@ 4.2 Qualitative Visual Improvements
Visual inspections of upscaled video frames reveal distinct strengths. Figure 15 shows the RRDB-based model’s outputs for a BVI-AOM sequence, with bicubic inputs yielding sharper edges and finer textures compared to bilinear inputs, particularly in static scenes. Also Figure 15 illustrates the Residual-based model’s outputs for a BVI-HOMTex sequence, where it maintains smoother transitions and better perceptual coherence, as reflected in its higher SSIM, despite slightly blurrier edges.
@@ 4.3 Comparison with Traditional (Singular) Image Upscaling Technologies
To validate the strength of the spatio-temporal approach in video super-resolution, the results are compared with the results obtained using one of the most well-known single-image super-resolution (SISR) model, Real-ESRGAN. SISR methods, while effective for static images, often lack temporal coherence when applied frame-by-frame to videos, leading to flickering artifacts and reduced perceptual quality. The spatio-temporal SRGAN framework built in this thesis, incorporating 3D Non-Local Blocks and temporal feature aggregation, addresses these limitations by leveraging inter-frame dependencies, as evidenced by the performance on the BVI-AOM dataset. Further examples are shown in Figure  16.  [15]
@@ 4.3.2 Qualitative Comparison
SISR methods like Real-ESRGAN produce photo-realistic images but struggle with temporal consistency in video applications, resulting in flickering or inconsistent textures across frames, particularly on BVI-AOM (LPIPS 0.341, SSIM 0.865) and REDS (LPIPS 0.367, SSIM 0.875). The spatio-temporal RRDB-based model reduces such artifacts by leveraging temporal information, as seen in smoother transitions in BVI-AOM sequences (SSIM 0.973, LPIPS 0.140) (Figure 15). The RRDB-based model’s superior SSIM on REDS (0.927 vs. 0.875 for Real-ESRGAN) and BVI-HOMTex (0.924 vs. 0.875) suggests better handling of motion and texture preservation, a critical advantage over SISR methods that treat frames independently. For comparison, iSeeBetter, another spatio-temporal VSR methods achieve a PSNR of 28.20 at most on the Vid4 dataset with GAN-based approaches on 4x upscaling task. On the other side, using only residual blocks with a simple non-local block, the Residual-based model’s SSIM and PSNR values reached to 0.90864 and 30.0833, respectively (on 2x upscaling task).

### v2

@@ 1.1 Background
Figure 1: Basic overview of a GAN architecture  [13]
@@
Figure 2: Basic overview of an SRGAN architecture  [13]
@@
Also in the following year in 2018, another architecture called as ESRGAN is being publicated and shown the literature that further upgrades can still be made on residual blocks by stacking them consecutively and let them share information among each other. Each output of those blocks are stored and shared directly with the any other residual block waiting on the flow way. The desired architecture is shown in Fig. (9) and called as Residual in Residual Dense Block (RRDB)  [16].
@@
Figure 5: a. Leaky ReLU compared to b. (ordinary) ReLU  [18]
@@
Figure 6: Spacetime View of Non-Local Block  [5]
@@ 2.1 Spatio-Temporal Learning Architecture built for Video Super-Resolution
At first, it should be noted that two different experimental and basic models for different complexities of jobs are designed. The first shown in Fig. (8) is built using consecutive RRDBs mainly, whereas the other model which is shown in Fig. (8) is built based on successive residual blocks. Both of the models can be split into 4 different parts, that they can be called channel expanding and temporal feature extraction layer, spatial feature extraction layer, reconstruction layer and upsampling and channel compression layer. The RRDBs are based on Residual Blocks, and each residual block is simply a sequentially connected series of convolutions with skip connections, where flow diagram of those blocks are shown in Fig. (9).a and shown in Fig. (9).b. In simple terms, each RRDB is also a series (3) of residual blocks with another skip connection at the end.
@@
Figure 9: a. Residual Block and b. Residual in Residual Dense Block modules designed for this project.
@@
Assuming $\sigma(y)=LeakyReLU(y)$, a convolutional operation followed by a leaky relu function in the flow diagram mathematically represents the following Eq. (5).
@@
In the end, output of the residual block used in this project is as in given in Eq. (6) below.
@@
Note the following notations represent the same thing, given in Eq. (7) below.
@@
Correspondingly, RRDB class designed for this project is as follows given in Eq. (8) below.
@@
On the other hand, Non-Local Blocks has a different calculation logic from which its spatial feature extraction ability comes from. Normally, ordinary 2D convolution blocks and its derivations are based on extracting the temporal relationship, which limits the model’s learning ability to 2D dataset only. However, with non-local blocks, even in its simplest form, which is called dot product based non-local blocks, we get the ability to learn the 3D relationships by calculating the weighted sum of the 3D input sets. Mathematically, it is represented in a very generic form given in Eq. (9)  [5].
@@
where the output of a non-local block indeed is as follows in Eq. (10)
@@
  • $g(\cdot)$: It is the unary function g (given in Eq. (12)) computes a representation of the input signal at the position $j$.
@@
The normalization factor $C(\cdot)$ is given below in Eq. (11).
@@
The unary function $g(\cdot)$ is given below in Eq. (9).
@@
Pairwise function $f(\cdot)$ used in this function (dot product) is given below in Eq. (13).
@@
    $f(x_{i},x_{j})=\sigma(w_{f}^{T}\cdot\theta(x_{i})^{T}\mathbin{\|}\phi(x_{j}))$ (19)
    Please note that $\sigma(\cdot)=ReLU(\cdot)$.
Figure 10: U-Net architecture used as discriminator  [36]
In this adaptation process, the purpose of the discriminator shall not be changed deeply. Since their performance on image segmentation is proven and it is shown that their encoder & decoder based architecture is reliably well-performing along large scales of projects  [37], a U-Net architecture (see Fig. (10)) is built as the discriminator and it can be kept the same as long as it does well in distinguishing artificially made images from real images at first. After a while, it will converge to a non-zero fixed amount of error. Attempting to minimize this error excessively, bring it very close to zero, and significantly increase the complexity of the discriminator block in the process will negatively affect the overall procedure. The backpropagation of a non-zero loss from the discriminator block to the generator is necessary for a healthy process.
@@ 3.1.7 Evaluation and Definition of Loss Functions
    Figure 11: a. Moore Neighborhood compared to b. Von Neumann Neighborhood  [46]
@@
    Figure 12: Evaluation of Laplacian Pyramid  [49]
@@ 4.2 Qualitative Visual Improvements
Visual inspections of upscaled video frames reveal distinct strengths. Fig. (13) shows the RRDB-based model’s outputs for a BVI-AOM sequence, with bicubic inputs yielding sharper edges and finer textures compared to bilinear inputs, particularly in static scenes. Also Fig. (13) illustrates the Residual-based model’s outputs for a BVI-HOMTex sequence, where it maintains smoother transitions and better perceptual coherence, as reflected in its higher SSIM, despite slightly blurrier edges.
@@ 4.3 Comparison with Traditional (Singular) Image Upscaling Technologies
To validate the strength of the spatio-temporal approach in video super-resolution, the results are compared with the results obtained using one of the most well-known single-image super-resolution (SISR) model, Real-ESRGAN. SISR methods, while effective for static images, often lack temporal coherence when applied frame-by-frame to videos, leading to flickering artifacts and reduced perceptual quality. The spatio-temporal SRGAN framework built in this thesis, incorporating 3D Non-Local Blocks and temporal feature aggregation, addresses these limitations by leveraging inter-frame dependencies, as evidenced by the performance on the BVI-AOM dataset. Further examples are shown in Fig. (14).  [16]
@@ 4.3.2 Qualitative Comparison
SISR methods like Real-ESRGAN produce photo-realistic images but struggle with temporal consistency in video applications, resulting in flickering or inconsistent textures across frames, particularly on BVI-AOM (LPIPS 0.341, SSIM 0.865) and REDS (LPIPS 0.367, SSIM 0.875). The spatio-temporal RRDB-based model reduces such artifacts by leveraging temporal information, as seen in smoother transitions in BVI-AOM sequences (SSIM 0.973, LPIPS 0.140) (Fig. (13)). The RRDB-based model’s superior SSIM on REDS (0.927 vs. 0.875 for Real-ESRGAN) and BVI-HOMTex (0.924 vs. 0.875) suggests better handling of motion and texture preservation, a critical advantage over SISR methods that treat frames independently. For comparison, iSeeBetter, another spatio-temporal VSR methods achieve a PSNR of 28.20 at most on the Vid4 dataset with GAN-based approaches on 4x upscaling task. On the other side, using only residual blocks with a simple non-local block, the Residual-based model’s SSIM and PSNR values reached to 0.90864 and 30.0833, respectively (on 2x upscaling task).